OpenStudy (anonymous):

\[\frac{ 2 }{ 8x^2 }=\frac{ 1 }{ 8x }\]

OpenStudy (kropot72):

Start by cross multiplying. What do you get?

OpenStudy (anonymous):

I think it's x = 2

OpenStudy (kropot72):

How did you get that answer?

OpenStudy (anonymous):

umm.. one second..

OpenStudy (anonymous):

I cross multiplied and got 8x^2 = 2 * 8x, which is, 8x^2 = 16x, then I divided it by 8 and got x=2

OpenStudy (anonymous):

oops meant divided it by 8x

OpenStudy (kropot72):

Good work! Your answer is correct.

OpenStudy (anonymous):

yay! Can you help me with another one?

OpenStudy (kropot72):

Go ahead with one more.

OpenStudy (anonymous):

okay, that's the last one.

OpenStudy (anonymous):

Using complete sentences, explain the steps to solve the equation using the LCD method. \[\frac{ 3 }{ x+2 } + \frac{ 1 }{ 2x }=\frac{ 4 }{ x+2 }\]

OpenStudy (anonymous):

I'm thinking it's like x = 1, but I don't think it's right...

OpenStudy (kropot72):

The LCD of the denominators on the left hand side is 2x(x + 2) Using this LCD the left hand side becomes \[\frac{6x+x+2}{2x(x+2)}=\frac{7x+2}{2x(x+2)}\] So the equation now becomes \[\frac{7x+2}{2x(x+2)}=\frac{4}{x+2}\] What do you get if you multiply both sides of the equation by (x + 2)?

OpenStudy (anonymous):

Well the x+2 would cancel out. Right?

OpenStudy (kropot72):

Good work! so we now get \[\frac{7x+2}{2x}=\frac{4}{1}\] Now cross multiply and what do you get?

OpenStudy (anonymous):

umm 7x + 2 * 1 = 4 * 2x??

OpenStudy (kropot72):

Correct 7x + 2 =8x Now subtract 7x from both sides and you have it!

OpenStudy (anonymous):

umm. x = 2?

OpenStudy (kropot72):

Your answer is correct! Well done :)

OpenStudy (anonymous):

YAY!!!!!!!!!!! THANK YOU SOOOOOOOO MUCH!!! YOU'RE LIKE MY HERO!! CUZ I'VE BEEN WORKING ON THIS TEST FOR LIKE ONE AND A HALF HOURS!!!! :DDDD

OpenStudy (kropot72):

You're welcome :)